Check whether or not the set of all symmetric matrix in Mn(R) form a real vector space with respect to the usual addition and scalar multiplication for Mn(R)


1
Expert's answer
2021-03-01T07:15:06-0500

By definition, a real vector space is a non-empty set (it is true for "M_n \lparen \Reals \rparen" ) equipped with operations of addition and multiplication with a real number which satisfy the following axioms:

  1. "\forall A, B \in M_n \lparen \Reals \rparen A + B = B + A" ( it is true for arbitrary matrices)
  2. "\forall A, B, C \in M_n \lparen \Reals \rparen (A + B) + C = A + (B + C)" (it is true for arbitrary matrices)
  3. "\exists \theta \in M_n (\Reals) : \forall A \in M_n (\Reals) A + \theta = \theta + A = A" (it is true because the zero matrix is a symmetric matrix)
  4. "\forall A \in M_n (\Reals) \exists (-A) \in M_n (\Reals) : \\ A + (-A) = (-A) + A = \theta" (it follows from the matrix transposition properties)
  5. "\exists I \in M_n (\Reals) : \forall A \in M_n (R) I \cdot A = A" (it is true because the identical matrix is a symmetric matrix)
  6. "\forall \alpha , \beta \in \Reals , \forall A \in M_n (\Reals) (\alpha \cdot \beta) \cdot A = \alpha \cdot (\beta \cdot A)" (it follows from the matrix multiplication with scalar properies)
  7. "\forall \alpha \in \Reals , \forall A, B \in M_n (\Reals) \alpha \cdot (A + B) = \\ = \alpha \cdot A + \alpha \cdot B" (it follows from the matrix multiplication with scalar properties)
  8. "\forall \alpha , \beta \in \Reals , \forall A \in M_n (\Reals) (\alpha + \beta) \cdot A = \\ = \alpha \cdot A + \beta \cdot A" (it follows from the matrix multiplication with scalar properties)

Thus, "M_n (\Reals)" is a real vector space.